### v1

@@ 4 Regridding the sky into a Cartesian domain
A further challenge for intensity mapping analyses is the transformation of observational data from its native coordinates on the sky, $(\mathrm{R.A.},\mathrm{Dec.},\nu)$, into a three-dimensional Cartesian cube (x, y, z coordinates) in comoving units of $h^{-1}\mathrm{Mpc}$. This step is essential because estimators for clustering statistics are currently most efficiently applied in Fourier space, where Fast Fourier Transforms (FFTs) require data sampled on a regular grid [50]. Remaining in configuration space and using two-point correlation functions is a viable alternative [86, 17] with some useful properties, such as a more natural way to handle complex survey masks. However, as intensity mapping surveys scale up in size, counting pixel pairs will get increasingly computationally expensive. One could also, in principle, compute Fourier modes directly from sky voxels (without regular grid FFTs), but again the enormous size of modern surveys renders this computationally infeasible, making Cartesian regridding a critical prerequisite for Fourier-based (e.g. power spectra) analysis.
Many recent pathfinder studies approximated the survey footprint as a rectangular box in comoving space, assigning comoving lengths to the map edges and then applying an FFT directly [152, 78]. However, this approximation introduces significant biases, with power spectrum estimates offset by more than $20\%$ across scales, and up to $30\%$ on the largest modes, even for surveys with relatively modest sky coverage. The root cause lies in the non-cubical geometry of real surveys: intensity mapping observations discretise the sky into voxels of unequal comoving volume, forming truncated conical footprints rather than rectangular boxes. Neglecting this geometry leads to systematic distortions in measured clustering, and this will worsen as surveys scale up in size.
To address this, MeerKLASS currently adopts the framework developed in Cunnington and Wolz [53], with a particle resampling framework that redistributes the voxel intensities of the input sky maps into a uniform 3D Cartesian grid (see the left panel of Figure 2 for a visual demonstration). By populating each sky voxel with sampling particles whose coordinates are transformed into comoving space, and then assigning these to regular grid cells, the method can avoid large-scale projection distortions.
@@
Whilst this resampling will sufficiently sample the voxel intensities into a new Cartiesian frame with minimal distortion, a statistical analysis of this regridded field alone will reveal significant biases relative to the true model. This is shown by the blue lines in the right panel of Figure 2. Here the estimated power spectra $\hat{P}_{\rm G}$ from the regridded intensity mapping simulations are compared to the corresponding input model $P_{\rm mod}$. Results are displayed for two different simulation versions, the “Cubic voxels” and “Fine channel”, discussed more shortly. The power spectra are averaged over 100 simulation realisations [see 53, for more details]. The vertical dashed red lines indicate the Nyquist frequency $k_{\rm Nyq}$ and $0.5\,k_{\rm Nyq}$ of the Cartesian grid.
The blue curves show a clear suppression of power at high-$k$. This is caused by a smoothing to the field which is introduced by the nearest-grid-point (NGP) mass assignment of the resampling particles. Fortunately, this is a well-studied related problem in discrete galaxy catalogue-based clustering analyses [82, 122]. To mitigate this, one can correct for the effective convolution of the field with the mass assignment window function, $W_{\rm mas}(\bm{x})$. This is most simply done in Fourier space by dividing the field through by the Fourier-space transform of the assignment function,
@@
For NGP, effectively a top-hat in configuartion space, this corresponds to a product of sinc functions for each dimension in Fourier space [76]
@@
The difference between the Fine channel simulations is that the spacing in frequency along the line-of-sight is much finer, more resembling a typical radio intensity mapping survey. In the Fine channel case (dashed lines), the increased frequency resolution suppresses aliasing in the radial modes, leading to smaller deviations from the true model compared with the Cubic voxels case (solid lines). Figure 2 therefore demonstrates the two main systematic effects introduced by the regridding process: damping of small-scale power by particle assignment, and artificial enhancement by aliasing at high-$k$, the latter highlighting the need for further mitigation. This is also a problem shared by galaxy clustering analyses, and higher-order particle assignments and interlacing schemes exist [122], which can increase accuracy further if required. This was also explored in Cunnington and Wolz [53] and with targeted treatment, sub-percent accuracy could be achieved at up to 80% of the Nyquist frequency.
In current analyses [e.g. 100], we do not see evidence of aliasing due to the relatively low signal-to-noise from these early data sets; therefore, simple NGP assignments with no interlacing are adopted for simplicity, but the smoothing correction to the regridded field from particle sampling (Equation 11) is implemented. However, for future MeerKLASS observations, which are beginning to cover large sky areas with high sensitivity, adopting robust regridding methods is crucial to avoid spurious power spectrum features and to unlock the full cosmological potential of the survey.
@@ 4.1 Analysis beyond the flat-sky approximation
The current approach of regridding does not fundamentally escape the problem of a flat-sky approximation. In regridded fields, we must assume a global line-of-sight direction, and for a wide-field survey, this will not be mutually parallel with all lines-of-sight. To date, even the most advanced galaxy surveys [see e.g. 64] have relied on a flat-sky (or plane-parallel wave) approximation, embedding survey footprints into a Cartesian cube and applying ad-hoc wide-angle corrections. While this approach is workable for modest survey areas, its accuracy breaks down as coverage extends to wider areas, with biases creeping in for parameter inference (e.g. $f_{\rm NL}$) on the largest scales [26, 74]. For intensity mapping, the shortcomings are magnified because of the strongly anisotropic observational effects: frequency-dependent telescope beams that convolve purely angular modes, and foreground cleaning that preferentially removes large-scale radial modes. Accurately modelling or correcting these within a plane-parallel box will demand complex forward modelling and risks residual biases that undermine precision cosmology. A wide-angle correction recipe that also includes beam and similar configuration-space convolution effects is not currently available, but represents a viable option, whose effectiveness is currently being investigated within the MeerKLASS Collaboration.
@@
One alternative that circumvents many of these issues is to analyse the data in harmonic space, measuring angular power spectra $C_{\ell}$ as a function of frequency or redshift [16, 39]. This formalism naturally respects the curvature of the sky, avoids plane-parallel assumptions, and has been widely used with great success [11, 138, 139, 104, 69, 156]. Applied to intensity mapping, it would provide a clean framework to handle beam convolution and sky masks. However, harmonic analyses are intrinsically limited to angular clustering, effectively projecting out radial information within each frequency slice, unless a large number of radial bins are used, which then pushes the limits of computational feasibility [39, 124]. As a result, the exquisite tomographic potential of intensity mapping, its ability to probe many frequencies across a wide redshift range, is not fully exploited. Similarly, line-of-sight-only $P(\nu)$ power spectra are feasible [e.g. 143] and avoid regridding and plane-parallel approximations, but will face similar limitations, and will not fully exploit the full volume. This loss of three-dimensional information in $C_{\ell}$ and $P(\nu)$ approaches is particularly restrictive for ultra-large-scale cosmology, such as tests of primordial non-Gaussianity, which are optimal when they access the full radial and angular mode spectrum over vast cosmic volumes.
@@
Another intriguing solution is to use a fully 3D spherical treatment of the data. This is possible with the spherical Fourier–Bessel formalism [95, 42], which provides a native decomposition into angular and radial modes, automatically accommodating wide-angle terms and respecting the true geometry of the observations. Importantly, it also allows observational operators, such as the chromatic beam or foreground cleaning filters, to be applied where they physically act, either on angular or radial modes, avoiding the distortions introduced by Cartesian embedding. Early theoretical work has already highlighted the potential of SFB methods to recover large-scale clustering unbiasedly [148], plus there is potential for improving the computational feasibility of this approach [146, 113]. However, a robust end-to-end pipeline tailored to 21cm intensity mapping has yet to be developed. This is far beyond the scope of this work, but it is a planned exploration within the MeerKLASS collaboration.
@@ 5 Signal detections
The 21cm line of Hi offers a uniquely powerful probe of cosmology across an extensive redshift range, from the relatively recent Universe deep into the epoch of reionisation (EoR) [155, 88]. Much of the focus from the 21cm intensity mapping community is dedicated to this high-redshift frontier, aiming to detect the imprint of reionisation on the distribution of Hi. However, such measurements are extremely challenging: foregrounds are even more overwhelming at these low frequencies, and despite progressive improvements on upper limits using instruments such as LOFAR, MWA, and HERA [102, 105, 3], a full detection is yet to be obtained.
By contrast, one of the major strengths of 21cm intensity mapping at lower redshifts, in the post-EoR Universe, is the ability to probe the large-scale distribution of matter in synergy with other cosmological tracers. In particular, cross-correlation with optical galaxy surveys provides a robust pathway to detection: foregrounds and systematics in the radio maps are uncorrelated with optical data, while the cosmological signal is shared between the two [150, 30, 118]. This makes cross-correlation a powerful validation tool as well as a means to extract the first cosmological information from current-generation surveys, and characterise additive systematics.
In this section, we review some of the first detections of cosmological clustering in cross-correlation between early MeerKLASS intensity maps and overlapping galaxy surveys. These results represent a major milestone, demonstrating the feasibility of Hi intensity mapping for cosmology and paving the way for future measurements with the full MeerKLASS and, ultimately, the SKAO [128].
@@ 5.1 Cross-correlation power spectra with optical galaxies
The first detection of Hi intensity mapping with MeerKAT was reported in Cunnington et al. [56] using cross-correlation with galaxies from the WiggleZ Dark Energy Survey. With just 10.5 hours of L-band single-dish observations from the MeerKAT pilot survey covering  ${\sim}\,200\,$deg2 in the WiggleZ 11hr field, a $7.7\sigma$ detection was obtained of the cross-correlation power spectrum. This provided a measurement of the combined parameter $\Omega_{\textrm{H{i}}}b_{\textrm{H{i}}}r\,{=}\,[0.86\,{\pm}\,0.10\,{\rm(stat)}\,{\pm}\,0.12\,{\rm(sys)}]\,{\times}\,10^{-3}$ at an effective scale of $k_{\rm eff}\,{\sim}\,0.13\,h\,{\rm Mpc}^{-1}$. Crucially, this result represented the first practical demonstration of the multi-dish auto-correlation intensity mapping technique for cosmology, establishing MeerKAT as a pathfinder for SKAO-era science.
@@
Building on this, the MeerKLASS collaboration reported results from the L-band deep-field observations in MeerKLASS Collaboration et al. [100]. These maps, produced from 41 repeated scans over 236 deg2 (totalling 62 hours per dish before flagging), yielded the deepest (in terms of time integration) single-dish Hi intensity maps to date. Cross-correlating with a smaller overlapping spectroscopic sample from the GAMA survey, only covering about one third of the field, a detection of the cross-power spectrum was obtained at > $4\sigma$ significance in the redshift range 0.39<z< 0.460.39\,{<}\,z\,{<}\,0.46. This work showcased substantial improvements in calibration, building upon the work in Wang et al. [145] by including an iterative self-calibration. This work oversaw the transition into a regime where the thermal noise is no longer dominant compared to cosmological Hi fluctuations, necessitating more sophisticated covariance estimation strategies, which we discuss in the next sub-section.
@@
Most recently the work in Carucci et al. [41], revisited the pilot survey data overlapping with WiggleZ but focused on advancing the contaminant separation strategy. Introducing and testing a novel multiscale foreground-cleaning approach (mPCA, introduced in Section 3.1), this work demonstrated how a more efficient foreground cleaning can be more robust to signal loss. Without signal loss construction, this pipeline was able to confirm the cross-correlation detection at  ${\sim}\,6\sigma$ confidence, measuring $\Omega_{\textrm{H{i}}}b_{\textrm{H{i}}}r\,{=}\,[0.93\,{\pm}\,0.17]\,{\times}\,10^{-3}$ robustly across a wide range of scales ($0.04\,{\lesssim}\,k\,{\lesssim}\,0.3\,h\,{\rm Mpc}^{-1}$). This result instils greater confidence that MeerKAT Hi intensity mapping analyses can deliver unbiased cosmological information, laying further groundwork towards the full MeerKLASS.
@@
Figure 3 shows the three published cross-correlation power spectra overlaid on the same axes. The results are in good agreement within the signal-to-noise of the individual measurements, providing a consistent picture of the underlying Hi clustering. The shorter range of scales probed in the GAMA cross-correlation [100] is a consequence of the much smaller survey area of GAMA relative to WiggleZ, which limits the accessible Fourier modes. The extended reach to smaller $k$ achieved by Carucci et al. [41] arises from the more controlled multiscale foreground cleaning, which better preserves large-scale modes. All of these measurements were performed in the MeerKAT L-band, where RFI confines the usable redshift range to $0.39\,{\lesssim}\,z\,{\lesssim}\,0.46$. Future observations in the UHF band (outlined in Section 2) will dramatically expand the accessible redshift volume, allowing for much higher signal-to-noise detections and substantially improved cosmological constraints, which we discuss in more detail later in Section 6.
@@
Figure 3: Cross-correlation power spectra between MeerKAT L-band intensity maps and optical galaxy surveys at $z\,{\sim}\,0.43$. Blue-circles show the initial pilot survey detection with WiggleZ galaxies [56] (hollow markers indicate negative power), orange-diamonds the re-analysis of this data with improved foreground cleaning [41], and the green squares is the deep-field cross-correlation with GAMA [100]. Error bars show $1\sigma$ uncertainties estimated using different methods in each detection. The dashed lines are the corresponding model matched by colour. We refer the interested reader to the dedicated papers for details on these subtle distinctions.

### v2

@@ 4 Regridding the sky into a Cartesian domain
A further challenge for intensity mapping analyses is the transformation of observational data from its native coordinates on the sky, $(\mathrm{R.A.},\mathrm{Dec.},\nu)$, into a three-dimensional Cartesian cube (x, y, z coordinates) in comoving units of $h^{-1}\mathrm{Mpc}$. This step is essential because estimators for clustering statistics are currently most efficiently applied in Fourier space, where Fast Fourier Transforms (FFTs) require data sampled on a regular grid [Cooley1965AnAF]. Remaining in configuration space and using two-point correlation functions is a viable alternative [Kennedy:2021srz, Avila:2021wih] with some useful properties, such as a more natural way to handle complex survey masks. However, as intensity mapping surveys scale up in size, counting pixel pairs will get increasingly computationally expensive. One could also, in principle, compute Fourier modes directly from sky voxels (without regular grid FFTs), but again the enormous size of modern surveys renders this computationally infeasible, making Cartesian regridding a critical prerequisite for Fourier-based (e.g. power spectra) analysis.
Many recent pathfinder studies approximated the survey footprint as a rectangular box in comoving space, assigning comoving lengths to the map edges and then applying an FFT directly [Wolz:2015lwa, COMAP:2021sqw]. However, this approximation introduces significant biases, with power spectrum estimates offset by more than $20\%$ across scales, and up to $30\%$ on the largest modes, even for surveys with relatively modest sky coverage. The root cause lies in the non-cubical geometry of real surveys: intensity mapping observations discretise the sky into voxels of unequal comoving volume, forming truncated conical footprints rather than rectangular boxes. Neglecting this geometry leads to systematic distortions in measured clustering, and this will worsen as surveys scale up in size.
To address this, MeerKLASS currently adopts the framework developed in Cunnington:2023aou, with a particle resampling framework that redistributes the voxel intensities of the input sky maps into a uniform 3D Cartesian grid (see the left panel of Figure 2 for a visual demonstration). By populating each sky voxel with sampling particles whose coordinates are transformed into comoving space, and then assigning these to regular grid cells, the method can avoid large-scale projection distortions.
@@
Whilst this resampling will sufficiently sample the voxel intensities into a new Cartesian frame with minimal distortion, a statistical analysis of this regridded field alone will reveal significant biases relative to the true model. This is shown by the blue lines in the right panel of Figure 2. Here the estimated power spectra $\hat{P}_{\rm G}$ from the regridded intensity mapping simulations are compared to the corresponding input model $P_{\rm mod}$. Results are displayed for two different simulation versions, the “Cubic voxels” and “Fine channel”, discussed more shortly. The power spectra are averaged over 100 simulation realisations [see Cunnington:2023aou, for more details]. The vertical dashed red lines indicate the Nyquist frequency $k_{\rm Nyq}$ and $0.5\,k_{\rm Nyq}$ of the Cartesian grid.
The blue curves show a clear suppression of power at high-$k$. This is caused by a smoothing to the field which is introduced by the nearest-grid-point (NGP) mass assignment of the resampling particles. Fortunately, this is a well-studied related problem in discrete galaxy catalogue-based clustering analyses [Jing:2004fq, Sefusatti:2015aex]. To mitigate this, one can correct for the effective convolution of the field with the mass assignment window function, $W_{\rm mas}(\bm{x})$. This is most simply done in Fourier space by dividing the field through by the Fourier-space transform of the assignment function,
@@
For NGP, effectively a top-hat in configuartion space, this corresponds to a product of sinc functions for each dimension in Fourier space [HockneyEastwood]
@@
The difference between the Fine channel simulations is that the spacing in frequency along the line-of-sight is much finer, more resembling a typical radio intensity mapping survey. In the Fine channel case (dashed lines), the increased frequency resolution suppresses aliasing in the radial modes, leading to smaller deviations from the true model compared with the Cubic voxels case (solid lines). Figure 2 therefore demonstrates the two main systematic effects introduced by the regridding process: damping of small-scale power by particle assignment, and artificial enhancement by aliasing at high-$k$, the latter highlighting the need for further mitigation. This is also a problem shared by galaxy clustering analyses, and higher-order particle assignments and interlacing schemes exist [Sefusatti:2015aex], which can increase accuracy further if required. This was also explored in Cunnington:2023aou and with targeted treatment, sub-percent accuracy could be achieved at up to 80% of the Nyquist frequency.
In current analyses [e.g. MeerKLASS:2024ypg], we do not see evidence of aliasing due to the relatively low signal-to-noise from these early data sets; therefore, simple NGP assignments with no interlacing are adopted for simplicity, but the smoothing correction to the regridded field from particle sampling (Equation 11) is implemented. However, for future MeerKLASS observations, which are beginning to cover large sky areas with high sensitivity, adopting robust regridding methods is crucial to avoid spurious power spectrum features and to unlock the full cosmological potential of the survey.
@@ 4.1 Analysis beyond the flat-sky approximation
The current approach of regridding does not fundamentally escape the problem of a flat-sky approximation. In regridded fields, we must assume a global line-of-sight direction, and for a wide-field survey, this will not be mutually parallel with all lines-of-sight. To date, even the most advanced galaxy surveys [see e.g. DESI:2025zgx] have relied on a flat-sky (or plane-parallel wave) approximation, embedding survey footprints into a Cartesian cube and applying ad-hoc wide-angle corrections. While this approach is workable for modest survey areas, its accuracy breaks down as coverage extends to wider areas, with biases creeping in for parameter inference (e.g. $f_{\rm NL}$) on the largest scales [Benabou:2024tmn, 2025JCAP...07..063G]. For intensity mapping, the shortcomings are magnified because of the strongly anisotropic observational effects: frequency-dependent telescope beams that convolve purely angular modes, and foreground cleaning that preferentially removes large-scale radial modes. Accurately modelling or correcting these within a plane-parallel box will demand complex forward modelling and risks residual biases that undermine precision cosmology. A wide-angle correction recipe that also includes beam and similar configuration-space convolution effects is not currently available, but represents a viable option, whose effectiveness is currently being investigated within the MeerKLASS Collaboration.
@@
One alternative that circumvents many of these issues is to analyse the data in harmonic space, measuring angular power spectra $C_{\ell}$ as a function of frequency or redshift [Asorey:2012rd, Camera:2018jys]. This formalism naturally respects the curvature of the sky, avoids plane-parallel assumptions, and has been widely used with great success [Alonso:2018jzx, Tanidis:2020byi, Tanidis:2021uxp, Nakoneczny:2023nlt, DES:2024oud, Zheng:2025lqz]. Applied to intensity mapping, it would provide a clean framework to handle beam convolution and sky masks. However, harmonic analyses are intrinsically limited to angular clustering, effectively projecting out radial information within each frequency slice, unless a large number of radial bins are used, which then pushes the limits of computational feasibility [Camera:2018jys, Semenzato:2024rlc]. As a result, the exquisite tomographic potential of intensity mapping, its ability to probe many frequencies across a wide redshift range, is not fully exploited. Similarly, line-of-sight-only $P(\nu)$ power spectra are feasible [e.g. Villaescusa-Navarro:2016kbz] and avoid regridding and plane-parallel approximations, but will face similar limitations, and will not fully exploit the full volume. This loss of three-dimensional information in $C_{\ell}$ and $P(\nu)$ approaches is particularly restrictive for ultra-large-scale cosmology, such as tests of primordial non-Gaussianity, which are optimal when they access the full radial and angular mode spectrum over vast cosmic volumes.
@@
Another intriguing solution is to use a fully 3D spherical treatment of the data. This is possible with the spherical Fourier–Bessel formalism [Liu:2016xzv, Castorina:2017inr], which provides a native decomposition into angular and radial modes, automatically accommodating wide-angle terms and respecting the true geometry of the observations. Importantly, it also allows observational operators, such as the chromatic beam or foreground cleaning filters, to be applied where they physically act, either on angular or radial modes, avoiding the distortions introduced by Cartesian embedding. Early theoretical work has already highlighted the potential of SFB methods to recover large-scale clustering unbiasedly [Wen:2025zbv], plus there is potential for improving the computational feasibility of this approach [Wang:2020wsx, Raccanelli:2023zkj]. However, a robust end-to-end pipeline tailored to 21cm intensity mapping has yet to be developed. This is far beyond the scope of this work, but it is a planned exploration within the MeerKLASS collaboration.
@@ 5 Signal detections
The 21cm line of Hi offers a uniquely powerful probe of cosmology across an extensive redshift range, from the relatively recent Universe deep into the epoch of reionisation (EoR) [Zaroubi:2012in, Kovetz:2017agg]. Much of the focus from the 21cm intensity mapping community is dedicated to this high-redshift frontier, aiming to detect the imprint of reionisation on the distribution of Hi. However, such measurements are extremely challenging: foregrounds are even more overwhelming at these low frequencies (hundreds of times brighter at $\nu\,{<}\,200\,$MHz compared to 1,0001{,}000{\,}MHz), and despite progressive improvements on upper limits using instruments such as LOFAR, MWA, and HERA [Mertens:2025pvk, Nunhokee:2025jbn, HERA:2022wmy], a full detection is yet to be obtained.
By contrast, one of the major strengths of 21cm intensity mapping at lower redshifts, in the post-EoR Universe, is the ability to probe the large-scale distribution of matter in synergy with other cosmological tracers. In particular, cross-correlation with optical galaxy surveys provides a robust pathway to detection: foregrounds and systematics in the radio maps are uncorrelated with optical data, while the cosmological signal is shared between the two [Wolz:2015ckn, Berti:2023viz, Santos:2025fff]. This makes cross-correlation a powerful validation tool as well as a means to extract the first cosmological information from current-generation surveys, and characterise additive systematics.
In this section, we review some of the first detections of cosmological clustering in cross-correlation between early MeerKLASS intensity maps and overlapping galaxy surveys. These results represent a major milestone, demonstrating the feasibility of Hi intensity mapping for cosmology and paving the way for future measurements with the full MeerKLASS and, ultimately, the SKAO [SKA:2018ckk].
@@ 5.1 Cross-correlation power spectra with optical galaxies
The first detection of Hi intensity mapping with MeerKAT was reported in Cunnington:2022uzo using cross-correlation with galaxies from the WiggleZ Dark Energy Survey. With just 10.5 hours of L-band single-dish observations from the MeerKAT pilot survey covering  ${\sim}\,200\,$deg2 in the WiggleZ 11hr field, a $7.7\sigma$ detection was obtained of the cross-correlation power spectrum. This provided a measurement of the combined parameter $\Omega_{\textrm{H{i}}}b_{\textrm{H{i}}}r\,{=}\,[0.86\,{\pm}\,0.10\,{\rm(stat)}\,{\pm}\,0.12\,{\rm(sys)}]\,{\times}\,10^{-3}$ at an effective scale of $k_{\rm eff}\,{\sim}\,0.13\,h\,{\rm Mpc}^{-1}$. Crucially, this result represented the first practical demonstration of the the single-dish mode employed on a multi-dish instrument for cosmological intensity mapping, establishing MeerKAT as a pathfinder for SKAO-era science.
@@
Building on this, the MeerKLASS collaboration reported results from the L-band deep-field observations in MeerKLASS:2024ypg. These maps, produced from 41 repeated scans over 236 deg2 (totalling 62 hours per dish before flagging), yielded the deepest (in terms of time integration) single-dish Hi intensity maps to date. Cross-correlating with a smaller overlapping spectroscopic sample from the GAMA survey, only covering about one third of the field, a detection of the cross-power spectrum was obtained at > $4\sigma$ significance in the redshift range 0.39<z< 0.460.39\,{<}\,z\,{<}\,0.46. This work showcased substantial improvements in calibration, building upon the work in Wang:2020lkn by including an iterative self-calibration. This work oversaw the transition into a regime where the thermal noise is no longer dominant compared to cosmological Hi fluctuations, necessitating more sophisticated covariance estimation strategies, which we discuss in the next sub-section.
@@
Most recently the work in Carucci:2024qpm, revisited the pilot survey data overlapping with WiggleZ but focused on advancing the contaminant separation strategy. Introducing and testing a novel multiscale foreground-cleaning approach (mPCA, introduced in Section 3.1), this work demonstrated how a more efficient foreground cleaning can be more robust to signal loss. Without signal loss reconstruction, this pipeline was able to confirm the cross-correlation detection at  ${\sim}\,6\sigma$ confidence, measuring $\Omega_{\textrm{H{i}}}b_{\textrm{H{i}}}r\,{=}\,[0.93\,{\pm}\,0.17]\,{\times}\,10^{-3}$ robustly across a wide range of scales ($0.04\,{\lesssim}\,k\,{\lesssim}\,0.3\,h\,{\rm Mpc}^{-1}$). This result instils greater confidence that MeerKAT Hi intensity mapping analyses can deliver unbiased cosmological information, laying further groundwork towards the full MeerKLASS.
@@
Figure 3 shows the three published cross-correlation power spectra overlaid on the same axes. The results are in good agreement within the signal-to-noise of the individual measurements, providing a consistent picture of the underlying Hi clustering. The smaller range of scales probed in the GAMA cross-correlation [MeerKLASS:2024ypg] is a consequence of the much smaller survey area of GAMA relative to WiggleZ, which limits the accessible Fourier modes. The extended reach to smaller $k$ achieved by Carucci:2024qpm arises from the more controlled multiscale foreground cleaning, which better preserves large-scale modes. All of these measurements were performed in the MeerKAT L-band, where RFI confines the usable redshift range to $0.39\,{\lesssim}\,z\,{\lesssim}\,0.46$. Future observations in the UHF band (outlined in Section 2) will dramatically expand the accessible redshift volume, allowing for much higher signal-to-noise detections and substantially improved cosmological constraints, which we discuss in more detail later in Section 6.
@@
Figure 3: Cross-correlation power spectra between MeerKAT L-band intensity maps and optical galaxy surveys at $z\,{\sim}\,0.43$. Blue-circles show the initial pilot survey detection with WiggleZ galaxies [Cunnington:2022uzo] (hollow markers indicate negative power), orange-diamonds the re-analysis of this data with improved foreground cleaning [Carucci:2024qpm], and the green squares is the deep-field cross-correlation with GAMA [MeerKLASS:2024ypg]. Error bars show $1\sigma$ uncertainties estimated using different methods in each detection. The dashed lines are the corresponding model matched by colour. We refer the interested reader to the dedicated papers for details on these subtle distinctions.